# Augmenting Subspace Optimization Methods with Linear Bandits

### v1

@@ 15 Return: $S,S^{\perp},Q$
1 (Initialization) Choose algorithmic constants $\eta_{1},\eta_{2},\Delta_{\max}>0$ and $0<\nu_{1}<1<\nu_{2}$.
2 Choose horizon $K$ and regularizer $\lambda>0$. Choose initial point $x^{0}\in\mathbb{R}^{d}$ and initial trust-region radius $\Delta_{0}\in(0,\Delta_{\max})$.
3 Initialize $C_{1}\leftarrow\lambda I_{d}$, $b_{1}\leftarrow 0\in\mathbb{R}^{d}$, gradient estimate $g_{1}\leftarrow 0\in\mathbb{R}^{d}$, memory parameter $M$, average weighting parameter $\mu\in(0,1)$, initial gradient upper bound estimate $U_{1}$. Initialize a bank of points $\mathcal{Y}$ with pairs $(x,f(x))$ for which $f(x)$ is known.
@@
5       (Get initial subspace) Use Algorithm 5 to obtain $S_{k}$, $S_{k}^{\perp}$ and $Q_{k}$.
6       if $S_{k}^{\perp}\neq\emptyset$ then
7             (Determine one-dimensional sketch from UCB) Find $s_{k}$ by solving (13), $S_{k}\leftarrow[S_{k},s_{k}]$.
@@
9      (Obtain additional random sketches) Choose $p_{k}\in[0,d-dim(span(S_{k}))]$ and call Algorithm 7 to obtain $Q_{k}\in\mathbb{R}^{d\times p_{k}}$, $S_{k}\leftarrow[S_{k},Q_{k}]$.
10       (Perform additional function evaluations) Evaluate $\{f(x^{k}+\Delta_{k}q_{i}):q_{i}\in s_{k}\cup Q_{k}\}$ and update $\mathcal{Y}$.
11       (Build model) Build a model $m_{k}:\mathbb{R}^{dim(span(S_{k}))}\to\mathbb{R}$ of $f$ valid on the affine subspace defined by $x_{k}$ and $S_{k}$.
12       (Solve TRSP) (Approximately) solve $\min_{z:\|z\|\leq\Delta_{k}}m_{k}(z)$ to obtain $z_{k}$.
13       (Evaluate new point) Evaluate $f(x_{k}+S_{k}z_{k})$ and update $\mathcal{Y}$.
14       (Determine acceptance) Compute $\rho_{k}\leftarrow\displaystyle\frac{f(x_{k})-f(x_{k}+S_{k}z_{k})}{m_{k}(0)-m_{k}(z^{k})}$.
@@
16             $x_{k+1}\leftarrow x_{k}+S_{k}z_{k}$.
@@
20             if $\|\nabla m_{k}(0)\|\geq\eta_{2}\Delta_{k}$ then
@@
23                  $\Delta_{k+1}\leftarrow\nu_{1}\Delta_{k}$
24
@@ 23                  $\Delta_{k+1}\leftarrow\nu_{1}\Delta_{k}$
27      (Update UCB estimators) Update $C_{k+1}$ and $b_{k+1}$ according to (12), replacing $r_{j,i}$ with corresponding entries of $\nabla m_{j}(0)$.
28       $g_{k+1}\leftarrow C_{k+1}^{-1}b_{k+1}$.
29       $U_{k+1}\leftarrow\mu U_{k}+(1-\mu)\frac{d}{dim(span(S_{k}))}\|\nabla m_{k}(0)\|.$
30
@@
1 Input: Orthogonal basis for subspace $S$, number of new directions $p_{k}$.
2 Randomly generate $A\in\mathbb{R}^{d\times p_{k}}$ from some distribution.
3 $A\leftarrow A-SS^{\top}A$
Perform the QR factorization of $A$, and return the first $p_{k}$ columns of the orthonormal $Q$ factor.
Algorithm 7 Generating orthogonal random sketches in $S^{\perp}$
@@ 5 Numerical Tests
The hypothesis motivating the experiments performed in this section is that, without any sophisticated hyperparameter tuning, employing a sliding window UCB mechanism in addition to a standard randomized generation of $S_{k}$ generally leads to practical gains over a method that performs only randomized sketches.
The work in this manuscript was motivated primarily by applications where gradients are computationally expensive to compute. Thus we implicitly assume that all linear algebra costs incurred in updating the (inverse) covariance matrix and the vector $b$, as well as the costs of performing multiplications with the inverse covariance matrix in solving (6), are dominated by the cost of computing a (sketched) gradient. Thus, to model this sort of computational setting, and for the sake of these experiments, we are interested solely in maximizing objective function decrease given a fixed budget of oracle accesses.
In our experiments, and motivated by our discussion of when subspace optimization methods within Framework 1 are most appropriate, we implemented and tested reasonably faithful versions of both Algorithm 4 and Algorithm 6. Algorithm 4, in both Algorithm 4 and Algorithm 4, assumes access to an implementation of forward-mode AD to return arbitrary directional derivatives. For easy reproducibility, however, we chose in our experiments to use a large collection of unconstrained problems from CUTEst [20]. Because gradients in CUTEst are computed analytically and are relatively inexpensive to obtain by design, we simulate a forward AD oracle by computing the full gradient $\nabla f(x_{k})$, but only allowing the solver to observe the sketched quantity $S_{k}^{\top}\nabla f(x_{k})$, given the solver’s choice of $S_{k}$. SS-POUNDers, our implementation of Algorithm 6, is derivative-free.
@@ 5.1 Testing Algorithm 4
With our hypothesis and computational setting fixed, we consider two versions of Algorithm 4. Both methods have access to an oracle that will compute a sketch of size $p$. The first version of the method, which we refer to simply as random-only, is purely randomized; in the $k$th iteration, the random-only method generates $S_{k}\in\mathbb{R}^{d\times p}$ as a Gaussian matrix. Algorithm 4 through Algorithm 4 of Algorithm 4, that is, all the lines that pertain to generating an upper confidence bound on the gradient, are omitted entirely in the random-only method. The second method, which we refer to as UCB, does not omit Algorithm 4 through Algorithm 4 of Algorithm 4. In order to give maximally equal oracle access to the two methods, the same sequence of random Gaussian matrices $S_{k}$ that the random-only method employs is also given to the UCB method but with the last column of $S_{k}$ removed. Thus, both methods receive the same number of one-dimensional sketches in every iteration from the oracle, and the randomized directional information they receive is comparable in that, in the $k$th iteration, $p-1$ of the directional derivatives computed are equivalent across both methods; the difference is that the random-only method effectively chooses a random Gaussian vector as its $p$th direction, while the UCB method solves eq. 6 to determine its $p$th direction.
@@
$r_{i,p}:=\displaystyle\frac{f(x^{rand,i,p})-f(x^{ucb,i,p})}{\max\{f(x_{0})-f(x^{rand,i,p}),f(x_{0})-f(x^{ucb,i,p}),1\}}\in[-1,1],$ (14)
where $i$ indexes the 30 seeds, $p$ indexes the problem set $\mathcal{P}$, $x^{rand,i,p}$ denotes the last incumbent of the random-only variant on the $i$th seed applied to problem $p$, $x^{ucb,i,p}$ denotes the last incumbent of the UCB variant on the $i$th seed applied to problem $p$, and $x_{0}$ denotes a common initial point. Each box-and-whisker plot in our figures corresponds to a single problem $p$ and illustrates the 30 values of $r_{i,p}$. From the definition of $r_{i,p}$ in (14), we see that the UCB variant performs better at the minimization task (1) on a single run with a fixed seed whenever $r_{i,p}>0$. We separate our problems by problem size, with the smallest problems (dimension $11\leq d\leq 100$) in Figure 1, slightly larger problems (dimension $101\leq d\leq 1000$) in Figure 2, and the largest problems (dimension $1001\leq d\leq 10000$) in Figure 3. In all our tests, the horizon was chosen as $K=1000$. For a high-dimensional problem with $d=10000$ and $p_{k}=\lceil 0.001d\rceil$, this amounts to a total of $10,000$ directional derivatives, that is, the same amount of effort that would be required to compute a single full-space gradient. Thus, these tests are meant to simulate a true computationally expensive setting that would preclude the use of full gradients. Our choice to employ relative performance ratios (14) in these tests is thus justified, since they represent relative performance within such a constrained budget.
@@
Figure 1: Values of $r_{i,p}$ (see (14)) for unconstrained CUTEst problems of dimension $11\leq d\leq 100$. Top: Results with $p_{k}=\lceil 0.1d\rceil$. Bottom: Results with $p_{k}=\lceil 0.01d\rceil$.
Figure 2: Values of $r_{i,p}$ (see (14)) for unconstrained CUTEst problems of dimension $101\leq d\leq 1000$. Top: Results with $p_{k}=\lceil 0.01d\rceil$. Bottom: Results with $p_{k}=\lceil 0.001d\rceil$.
Figure 3: Values of $r_{i,p}$ (see (14)) for unconstrained CUTEst problems of dimension $1001\leq d\leq 10000$. Top: Results with $p_{k}=\lceil 0.01d\rceil$. Bottom: Results with $p_{k}=\lceil 0.001d\rceil$.
At a high level, Figure 1, Figure 2 and Figure 3 illustrate what we had hoped to show—across all problem sizes, there is a preference for using UCB over random-only, in the sense that most ratios across all runs are greater than 0.0, with particular exceptions seen in one problem in the $101\leq d\leq 1000$ setting (PENALTY2) and one problem in the $1001\leq d\leq 10000$ setting (SSCOSINE). In general, we observe a trend that as the dimension $d$ increases, the utility of UCB increases. Moreover, for a fixed problem set of similar dimension, UCB is generally more effective when smaller sample sizes $p_{k}$ are employed.
@@ 5.2 Testing SS-POUNDers
We tested our implementation of SS-POUNDers, which borrows heavily from subroutines found in the Matlab implementation of POUNDers available in IBCDFO [47]. We first demonstrate the performance of SS-POUNDers on all 44 of the unconstrained YATSOp [1] problems labeled midscale (i.e., with $98\leq d\leq 125$). We note that the YATSOp problems are a Matlab implementation of a particular subset of nonlinear least squares problems from CUTEst [20]. Like the POUNDers software from which SS-POUNDers is derived, the “default” use case of SS-POUNDers is for such nonlinear least squares problems. This specification to nonlinear least squares essentially only affects how model gradients and model Hessians are computed from residual vector evaluations in Algorithm 6 of Algorithm 6; the geometry point selection determined by the UCB mechanism in Algorithm 6 of Algorithm 6 is agnostic to this additional nonlinear least squares structure.
@@
We tested three variants of SS-POUNDers for illustrative purposes. The first variant, which we simply denote “UCB” in the figure legends, performs Algorithm 6 of Algorithm 6 but effectively skips Algorithm 6 of Algorithm 6 by always setting $p_{k}=0$. Note that this UCB variant, like POUNDers, is a deterministic method. The second variant, which we denote “random-only” in the figure legends, skips Algorithm 6 of Algorithm 6 entirely but sets $p_{k}=\min\{d-dim(span(S_{k})),1\}$ in Algorithm 6 of Algorithm 6. This yields a fully randomized derivative-free subspace optimization method like several of those discussed in our literature review but specialized to the POUNDers framework. Our third variant, denoted “UCB + random” in the figure legends, performs Algorithm 6 of Algorithm 6 and also sets $p_{k}=\min\{d-dim(span(S_{k})),1\}$ in Algorithm 6 of Algorithm 6. This corresponds to the successful UCB variant of Algorithm 4 from Section 5.1.
We show comparisons in terms of data profiles [36]. Given a set of problems $\mathcal{P}$ (each problem $prob$ is assumed to have a common initial point $x^{prob}_{0}$) and a set of solvers $\mathcal{S}$, denote the sets of points evaluated by a solver $sol\in\mathcal{S}$ when solving an instance of problem $prob\in\mathcal{P}$, $\{x^{sol,prob}_{k}\}_{k=1}^{K}$. Given a tolerance $\tau\in(0,1)$, we say that a solver $sol\in\mathcal{S}$ solved a problem $prob\in\mathcal{P}$ to tolerance $\tau$ within $N\leq K$ iterations provided
$f(x^{sol,prob}_{N})\leq\tau\left[f(x^{prob}_{0})-\displaystyle\min_{sol^{^{\prime}}\in\mathcal{S}}\displaystyle\min_{k=0,1,\dots,K}f(x^{sol^{^{\prime}},prob}_{k})\right].$
A data profile then fixes a tolerance $\tau$ and generates a line for each solver $sol\in\mathcal{S}$. On the $x$-axis, a data profile illustrates the number of budget units expended (in this manuscript, these budget units will always be groups of $d+2$ many function evaluations); and on the $y$-axis, a data profile shows the percentage of problems solved by solver $sol$ to tolerance $\tau$ within the corresponding budget. The choice of $d+2$ is because, in the absence of prior function evaluations being provided, POUNDers requires $d+1$ function evaluations to construct an initial affine model and then requires one additional evaluation to evaluate the trial point that would be suggested by that initial model in the first iteration. That is, a function evaluation budget of $d+2$ guarantees POUNDers can complete one full iteration in the absence of prior evaluations. For each problem/solver combination, we ran 10 fixed seeds of each of the randomized methods (the “random-only” and “UCB + random” variants), and we treated each seed as its own problem instance.
@@
As in the experiments of Section 5.1, we first demonstrate results for a very budget-constrained computational setting, in particular, when only $d+2$ function evaluations—that is, one budget unit—is allotted. Unsurprisingly, in this extremely budget-constrained budget, and as illustrated in Figure 4, POUNDers performs poorly compared with all of the SS-POUNDers variants. Once the budget is increased to a level where POUNDers is able to identify a reasonable solution ($50*(d+2)$ function evaluations in these experiments), a preference for using POUNDers re-emerges, especially when we demand tighter tolerances of $\tau=0.01$ or $\tau=0.001$; see Figure 5. Remarkably, however, at these tolerances the variant of SS-POUNDers employing both the UCB mechanism and one-dimensional random sketches is not considerably worse than POUNDers in the latter setting. From this particular experiment, however, we do see a general preference for the UCB+random variant of SS-POUNDers over the other two variants of Algorithm 6. Importantly, we see no evidence that one should employ a fully randomized variant of SS-POUNDers over a variant that employs both a UCB mechanism and randomized sketching.
@@
Next we test a hypothesis made earlier in this manuscript that employing a UCB mechanism should exhibit stronger performance than a full-space method on functions exhibiting low effective dimensionality. For this test we use the low-dimensional ($2\leq d\leq 12$) Moré–Wild benchmarking set for DFO [36]. Following the experimental setup in [9] and originally proposed in [46], we specify a dimension $D$ (in the tests illustrated here, $D=100$) and artificially construct objective functions with $D$-dimensional domains starting from the benchmark problems. This is done by taking a function $f:\mathbb{R}^{d}\to\mathbb{R}$ from the benchmark set and defining a new function $\bar{f}:\mathbb{R}^{D}\to\mathbb{R}$. Specifically, we first define an intermediate function $\tilde{f}:\mathbb{R}^{D}\to\mathbb{R}$ that acts trivially on the last $D-d$ dimensions of its domain, that is, $\tilde{f}(x)=f(\left[x_{1},\dots,x_{d}\right]^{\top})$. We then define $\bar{f}=\tilde{f}(Qx)$, where $Q\in\mathbb{R}^{D\times D}$ is a randomly generated orthogonal matrix; this rotation is done to prevent the subspace of low effective dimensionality from being always coordinate-aligned. Comparisons of the same variants of SS-POUNDers and the same experimental setup as in the prior set of experiments are shown in Figure 6. We see that in this setting of low effective dimensionality, subspace optimization methods are clearly superior to the full-space POUNDers. Notably, at all three tolerances shown, the data profiles suggest a slight preference for using the deterministic UCB variant. This observation is reasonable given that there is a constant subspace of variation that the UCB mechanism can learn, and so the subspace method is more likely to thrive when the balance of exploration vs exploitation is tipped to the direction of exploitation, since exploration in the orthogonal subspace defined by the last $D-d$ columns of the random matrix $Q$ will never align with $\nabla\bar{f}(x)$.

### v4

@@ 15 Return: $\mathbf{S},\mathbf{S}^{\perp},\mathbf{Q}$
1(Initialization) Choose algorithmic constants $\eta_{1},\eta_{2},\Delta_{\max}>0$ and $0<\nu_{1}<1<\nu_{2}$.
2 Choose horizon $K$ and regularizer $\lambda>0$. Choose initial point $\mathbf{x}^{0}\in\mathbb{R}^{d}$ and initial trust-region radius $\Delta_{0}\in(0,\Delta_{\max})$.
3 Initialize $\mathbf{C}_{1}\leftarrow\lambda\mathbf{I}_{d}$, $\mathbf{b}_{1}\leftarrow\mathbf{0}\in\mathbb{R}^{d}$, gradient estimate $\mathbf{g}_{1}\leftarrow\mathbf{0}\in\mathbb{R}^{d}$, memory parameter $M$, average weighting parameter $\mu\in(0,1)$, initial gradient upper bound estimate $U_{1}$. Initialize a bank of points $\mathcal{Y}$ with pairs $(\mathbf{x},f(\mathbf{x}))$ for which $f(\mathbf{x})$ is known.
@@
5 (Get initial subspace) Use Algorithm 6 to obtain $\mathbf{S}_{k}$, $\mathbf{S}_{k}^{\perp}$ and $\mathbf{Q}_{k}$.
6 if $\mathbf{S}_{k}^{\perp}\neq\emptyset$ then
7    (Determine one-dimensional sketch from UCB) Find $\mathbf{s}_{k}$ by solving (12), $\mathbf{S}_{k}\leftarrow[\mathbf{S}_{k},\mathbf{s}_{k}]$.
@@
9 else
10    $\mathbf{s}_{k}\leftarrow\emptyset.$
11 (Obtain additional random sketches) Choose $p_{k}\in[0,d-\mathrm{dim}(\mathrm{Span}(\mathbf{S}_{k}))]$ and call Algorithm 8 to obtain $\mathbf{Q}_{k}\in\mathbb{R}^{d\times p_{k}}$, $\mathbf{S}_{k}\leftarrow[\mathbf{S}_{k},\mathbf{Q}_{k}]$.
[…]
13 (Build model) Build a model $m_{k}:\mathbb{R}^{\mathrm{dim}(\mathrm{Span}(\mathbf{S}_{k}))}\to\mathbb{R}$ of $f$ valid on the affine subspace defined by $\mathbf{x}_{k}$ and $\mathbf{S}_{k}$.
14 (Solve TRSP) (Approximately) solve $\min_{\mathbf{z}:\|\mathbf{z}\|\leq\Delta_{k}}m_{k}(\mathbf{z})$ to obtain $\mathbf{z}_{k}$.
15 (Evaluate new point) Evaluate $f(\mathbf{x}_{k}+\mathbf{S}_{k}\mathbf{z}_{k})$ and update $\mathcal{Y}$.
16 (Determine acceptance) Compute $\rho_{k}\leftarrow\displaystyle\frac{f(\mathbf{x}_{k})-f(\mathbf{x}_{k}+\mathbf{S}_{k}\mathbf{z}_{k})}{m_{k}(\mathbf{0})-m_{k}(\mathbf{z}^{k})}$.
@@
18    $\mathbf{x}_{k+1}\leftarrow\mathbf{x}_{k}+\mathbf{S}_{k}\mathbf{z}_{k}$.
@@
22    if $\|\nabla m_{k}(\mathbf{0})\|\geq\eta_{2}\Delta_{k}$ then
@@
25       $\Delta_{k+1}\leftarrow\nu_{1}\Delta_{k}$.
26
@@
29 (Update UCB estimators) Update $\mathbf{C}_{k+1}$ and $\mathbf{b}_{k+1}$ according to (11), replacing $r_{j,i}$ with corresponding entries of $\nabla m_{j}(\mathbf{0})$.
30 $\mathbf{g}_{k+1}\leftarrow\mathbf{C}_{k+1}^{-1}\mathbf{b}_{k+1}$.
31 $U_{k+1}\leftarrow\mu U_{k}+(1-\mu)\frac{d}{\mathrm{dim}(\mathrm{Span}(\mathbf{S}_{k}))}\|\nabla m_{k}(\mathbf{0})\|.$
32
@@
1Input: Orthogonal basis for subspace $\mathbf{S}$, number of new directions $p_{k}$.
2 Randomly generate $\mathbf{A}\in\mathbb{R}^{d\times p_{k}}$ from some distribution.
3 $\mathbf{A}\leftarrow\mathbf{A}-\mathbf{S}\mathbf{S}^{\top}\mathbf{A}$
Perform the QR factorization of $\mathbf{A}$, and return the first $p_{k}$ columns of the orthonormal $\mathbf{Q}$ factor.
Algorithm 8 Generating orthogonal random sketches in $\mathbf{S}^{\perp}$
@@ 5 Numerical Tests
The hypothesis motivating the experiments performed in this section is that, without any sophisticated hyperparameter tuning, employing a sliding window UCB mechanism in addition to a standard randomized generation of $\mathbf{S}_{k}$ generally leads to practical gains over a method that performs only randomized sketches.
The work in this manuscript was motivated primarily by applications where gradients are computationally expensive to compute. Thus we implicitly assume that all linear algebra costs incurred in updating the (inverse) covariance matrix and the vector $\mathbf{b}$, as well as the costs of performing multiplications with the inverse covariance matrix in solving (5), are dominated by the cost of computing a (sketched) gradient. Thus, to model this sort of computational setting, and for the sake of these experiments, we are interested solely in maximizing objective function decrease given a fixed budget of oracle accesses.
In our experiments, and motivated by our discussion of when subspace optimization methods within Framework 1 are most appropriate, we implemented and tested reasonably faithful versions of both Algorithm 5 and Algorithm 7. Algorithm 5, in both Algorithm 5 and Algorithm 5, assumes access to an implementation of forward-mode AD to return arbitrary directional derivatives. For easy reproducibility, however, we chose in our experiments to use a large collection of unconstrained problems from CUTEst [23]. Because gradients in CUTEst are computed analytically and are relatively inexpensive to obtain by design, we simulate a forward AD oracle by computing the full gradient $\nabla f(\mathbf{x}_{k})$, but only allowing the solver to observe the sketched quantity $\mathbf{S}_{k}^{\top}\nabla f(\mathbf{x}_{k})$, given the solver’s choice of $\mathbf{S}_{k}$. SS-POUNDers, our implementation of Algorithm 7, is derivative-free.
@@ 5.1 Testing Algorithm 5
With our hypothesis and computational setting fixed, we consider two versions of Algorithm 5. Both methods have access to an oracle that will compute a sketch of size $p$. The first version of the method, which we refer to simply as random-only, is purely randomized; in the $k$th iteration, the random-only method generates $\mathbf{S}_{k}\in\mathbb{R}^{d\times p}$ as a Gaussian matrix. Algorithm 5 through Algorithm 5 of Algorithm 5, that is, all the lines that pertain to generating an upper confidence bound on the gradient, are omitted entirely in the random-only method. The second method, which we refer to as UCB, does not omit Algorithm 5 through Algorithm 5 of Algorithm 5. In order to give maximally equal oracle access to the two methods, the same sequence of random Gaussian matrices $\mathbf{S}_{k}$ that the random-only method employs is also given to the UCB method but with the last column of $\mathbf{S}_{k}$ removed. Thus, both methods receive the same number of one-dimensional sketches in every iteration from the oracle, and the randomized directional information they receive is comparable in that, in the $k$th iteration, p1p-1 of the directional derivatives computed are equivalent across both methods; the difference is that the random-only method effectively chooses a random Gaussian vector as its $p$th direction, while the UCB method solves eq. 5 to determine its $p$th direction.
@@
$r_{i,p}:=\displaystyle\frac{f(\mathbf{x}^{rand,i,p})-f(\mathbf{x}^{ucb,i,p})}{\max\{f(\mathbf{x}_{0})-f(\mathbf{x}^{rand,i,p}),f(\mathbf{x}_{0})-f(\mathbf{x}^{ucb,i,p}),1\}}\in[-1,1],$ (13)
where $i$ indexes the 30 seeds, $p$ indexes the problem set $\mathcal{P}$, $\mathbf{x}^{rand,i,p}$ denotes the last incumbent of the random-only variant on the $i$th seed applied to problem $p$, $\mathbf{x}^{ucb,i,p}$ denotes the last incumbent of the UCB variant on the $i$th seed applied to problem $p$, and $\mathbf{x}_{0}$ denotes a common initial point. Each box-and-whisker plot in our figures corresponds to a single problem $p$ and illustrates the 30 values of $r_{i,p}$. From the definition of $r_{i,p}$ in (13), we see that the UCB variant performs better at the minimization task (1) on a single run with a fixed seed whenever $r_{i,p}>0$. We separate our problems by problem size, with the smallest problems (dimension $11\leq d\leq 100$) in Figure 1, slightly larger problems (dimension $101\leq d\leq 1000$) in Figure 2, and the largest problems (dimension $1001\leq d\leq 10000$) in Figure 3. In all our tests, the horizon was chosen as $K=1000$. For a high-dimensional problem with $d=10000$ and $p_{k}=\lceil 0.001d\rceil$, this amounts to a total of $10,000$ directional derivatives, that is, the same amount of effort that would be required to compute a single full-space gradient. Thus, these tests are meant to simulate a true computationally expensive setting that would preclude the use of full gradients. Our choice to employ relative performance ratios (13) in these tests is thus justified, since they represent relative performance within such a constrained budget.
@@
Figure 1: Values of $r_{i,p}$ (see (13)) shown as box-whisker plots on 30 random replications of applying Algorithm 5 to CUTEst problems of dimension $11\leq d\leq 100$. Top: Results with $p_{k}=\lceil 0.1d\rceil$. Bottom: Results with $p_{k}=\lceil 0.01d\rceil$.
Figure 2: Values of $r_{i,p}$ (see (13)) shown as box-whisker plots on 30 random replications of applying Algorithm 5 to CUTEst problems of dimension $101\leq d\leq 1000$. Top: Results with $p_{k}=\lceil 0.01d\rceil$. Bottom: Results with $p_{k}=\lceil 0.001d\rceil$.
Figure 3: Values of $r_{i,p}$ (see (13)) shown as box-whisker plots on 30 random replications of applying Algorithm 5 to CUTEst problems of dimension $1001\leq d\leq 10000$. Top: Results with $p_{k}=\lceil 0.01d\rceil$. Bottom: Results with $p_{k}=\lceil 0.001d\rceil$.
At a high level, Figures 1, LABEL:, 2, LABEL: and 3 illustrate what we had hoped to show—across all problem sizes, there is a preference for using UCB over random-only, in the sense that most ratios across all runs are greater than 0.0, with particular exceptions seen in one problem in the $101\leq d\leq 1000$ setting (PENALTY2) and one problem in the $1001\leq d\leq 10000$ setting (SSCOSINE). In general, we observe a trend that as the dimension $d$ increases, the utility of UCB increases. Moreover, for a fixed problem set of similar dimension, UCB is generally more effective when smaller sample sizes $p_{k}$ are employed.
@@ 5.2 Testing SS-POUNDers
We tested our implementation of SS-POUNDers, which borrows heavily from subroutines found in the Matlab implementation of POUNDers available in IBCDFO [54]. We first demonstrate the performance of SS-POUNDers on all 44 of the unconstrained YATSOp [57] problems labeled midscale (i.e., with $98\leq d\leq 125$). We note that the YATSOp problems are a Matlab implementation of a particular subset of nonlinear least squares problems from CUTEst [23]. Like the POUNDers software from which SS-POUNDers is derived, the “default” use case of SS-POUNDers is for such nonlinear least squares problems. This specification to nonlinear least squares essentially only affects how model gradients and Hessians are computed from residual vector evaluations in Algorithm 7 of Algorithm 7; the geometry point selection determined by the UCB mechanism in Algorithm 7 of Algorithm 7 is agnostic to this additional nonlinear least squares structure.
@@
We tested three variants of SS-POUNDers for illustrative purposes. The first variant, which we simply denote “UCB” in the figure legends, performs Algorithm 7 of Algorithm 7 but effectively skips Algorithm 7 of Algorithm 7 by always setting $p_{k}=0$. Note that this UCB variant, like POUNDers, is a deterministic method. The second variant, which we denote “random-only” in the figure legends, skips Algorithm 7 of Algorithm 7 entirely but sets $p_{k}=\min\{d-\mathrm{dim}(\mathrm{Span}(\mathbf{S}_{k})),1\}$ in Algorithm 7 of Algorithm 7. This yields a fully randomized derivative-free subspace optimization method like several of those discussed in our literature review but specialized to the POUNDers framework. Our third variant, denoted “UCB + random” in the figure legends, performs Algorithm 7 of Algorithm 7 and also sets $p_{k}=\min\{d-\mathrm{dim}(\mathrm{Span}(\mathbf{S}_{k})),1\}$ in Algorithm 7 of Algorithm 7. This corresponds to the successful UCB variant of Algorithm 5 from Section 5.1.
We show comparisons in terms of data profiles [42]. Given a set of problems $\mathcal{P}$ (each problem probprob is assumed to have a common initial point $\mathbf{x}^{prob}_{0}$) and a set of solvers $\mathcal{S}$, denote the sets of points evaluated by a solver $sol\in\mathcal{S}$ when solving an instance of problem $prob\in\mathcal{P}$, $\{\mathbf{x}^{sol,prob}_{k}\}_{k=1}^{K}$. Given a tolerance $\tau\in(0,1)$, we say that a solver $sol\in\mathcal{S}$ solved a problem $prob\in\mathcal{P}$ to tolerance $\tau$ within $N\leq K$ iterations provided
$f(\mathbf{x}^{sol,prob}_{N})\leq\tau\left[f(\mathbf{x}^{prob}_{0})-\displaystyle\min_{sol^{^{\prime}}\in\mathcal{S}}\displaystyle\min_{k=0,1,\dots,K}f(\mathbf{x}^{sol^{^{\prime}},prob}_{k})\right].$
A data profile then fixes a tolerance $\tau$ and generates a line for each solver $sol\in\mathcal{S}$. On the $x$-axis, a data profile illustrates the number of budget units expended (in this manuscript, these budget units will always be groups of $d+2$ many function evaluations); and on the $y$-axis, a data profile shows the percentage of problems solved by solver solsol to tolerance $\tau$ within the corresponding budget. The choice of $d+2$ is because, in the absence of prior function evaluations being provided, POUNDers requires $d+1$ function evaluations to construct an initial affine model and then requires one additional evaluation to evaluate the trial point that would be suggested by that initial model in the first iteration. That is, a function evaluation budget of $d+2$ guarantees POUNDers can complete one full iteration in the absence of prior evaluations. For each problem/solver combination, we ran 10 fixed seeds of each of the randomized methods (the “random-only” and “UCB + random” variants), and we treated each seed as its own problem instance.
@@
As in the experiments of Section 5.1, we first demonstrate results for a very budget-constrained computational setting, in particular, when only $d+2$ function evaluations—that is, one budget unit—is allotted. Unsurprisingly, in this extremely budget-constrained budget, and as illustrated in Figure 4, POUNDers performs poorly compared with all of the SS-POUNDers variants. Once the budget is increased to a level where POUNDers is able to identify a reasonable solution ($50(d+2)$ function evaluations in these experiments), a preference for using POUNDers re-emerges, especially when we demand tighter tolerances of $\tau=0.01$ or $\tau=0.001$; see Figure 5. Remarkably, however, at these tolerances the variant of SS-POUNDers employing both the UCB mechanism and one-dimensional random sketches is not considerably worse than POUNDers in the latter setting. From this particular experiment, however, we do see a general preference for the UCB+random variant of SS-POUNDers over the other two variants of Algorithm 7. Importantly, we see no evidence that one should employ a fully randomized variant of SS-POUNDers over a variant that employs both a UCB mechanism and randomized sketching.
@@
Next we test a hypothesis made earlier in this manuscript that employing a UCB mechanism should exhibit stronger performance than a full-space method on functions exhibiting low effective dimensionality. For this test we use the low-dimensional ($2\leq~d\leq 12$) Moré–Wild benchmarking set for DFO [42]. Following the experimental setup in [9] and originally proposed in [53], we specify a dimension $D$ (in the tests illustrated here, $D=100$) and artificially construct objective functions with $D$-dimensional domains starting from the benchmark problems. This is done by taking a function $f:\mathbb{R}^{d}\to\mathbb{R}$ from the benchmark set and defining a new function $\bar{f}:\mathbb{R}^{D}\to\mathbb{R}$. Specifically, we first define an intermediate function $\tilde{f}:\mathbb{R}^{D}\to\mathbb{R}$ that acts trivially on the last DdD-d dimensions of its domain, that is, $\tilde{f}(\mathbf{x})=f(\left[\mathbf{x}_{1},\dots,\mathbf{x}_{d}\right]^{\top})$. We then define $\bar{f}=\tilde{f}(\mathbf{Q}\mathbf{x})$, where $\mathbf{Q}\in\mathbb{R}^{D\times D}$ is a randomly generated orthogonal matrix; this rotation is done to prevent the subspace of low effective dimensionality from being always coordinate-aligned. Comparisons of the same variants of SS-POUNDers and the same experimental setup as in the prior set of experiments are shown in Figure 6. We see that in this setting of low effective dimensionality, subspace optimization methods are clearly superior to the full-space POUNDers. Notably, at all three tolerances shown, the data profiles suggest a slight preference for using the deterministic UCB variant. This observation is reasonable given that there is a constant subspace of variation that the UCB mechanism can learn, and so the subspace method is more likely to thrive when the balance of exploration vs exploitation is tipped to the direction of exploitation, since exploration in the orthogonal subspace defined by the last DdD-d columns of the random matrix $\mathbf{Q}$ will never align with $\nabla\bar{f}(\mathbf{x})$.